OpenStudy (anonymous):

is y = c e^(ax) a) linear-linear b)log-linear c)log-log How about y = k/(x^1/2)??

OpenStudy (mani_jha):

From next time, please post such questions in the Mathematics section Take log on both sides \[\log y = \log (ce ^{ax})\] \[\log y = \log c + axlog e\] \[\log y = \log c + ax\] Now, tell me what does the above equation tell you about the answer?